# On Doubly Robust Estimation with Nonignorable Missing Data Using Instrumental Variables

### v2

@@ 2.3 Alternative factorization
$p(r,y,x;\varphi)=\pi(y,x;\eta,\gamma)^{r}\{1-\pi(y,x;\eta,\gamma)\}^{1-r}\underbrace{p(y\mid u;\psi)p(z\mid u;\beta)p(u;\zeta)}_{p(y,x;\psi,\beta,\zeta)},$
where $\gamma$ is finite-dimensional and $(\eta,\psi,\beta,\zeta)$ is infinite-dimensional. The full data density is clearly separated from the nonresponse mechanism in the factorization above. This naturally encodes the exclusion restriction of assumption 2, which operates on the full data distribution, rather than on subpopulations defined by the nonresponse status. Furthermore, $\psi$ and $\beta$ are variationally independent, in the sense that any appropriate choice of $\psi$ and $\beta$ would result in a density in $\mathcal{M}$. Therefore, the models $p(y\mid u;\psi)$ and $p(z\mid u;\beta)$ can be specified separately without concerns about incompatibility. We show in the paper that estimation of $\mu_{0}$ requires consistent estimation of at least a subset of the nuisance parameters $(\eta,\psi,\beta)$. In modern studies, a broad collection of baseline covariates and operational features are usually recorded. When $X$ is high-dimensional or contains several continuous components, nonparametric estimation is typically infeasible in the moderate sample sizes that are found in practice, as the data are too sparse due to the curse of dimensionality (Robins and Ritov,, 1997). Thus, we are often forced to specify more stringent dimension-reducing models $\eta(x;\xi)$, $p(y\mid u;\psi)$ and $p(z\mid u;\beta)$. Although in principle these models could be made as flexible as allowed by the sample size, we focus on parametric specifications in this paper to ease exposition.
@@
Accordingly, we define the submodels $\mathcal{M}_{1}$, $\mathcal{M}_{2}$ of $\mathcal{M}$ which correspond to models (1), (2) respectively. Thus, the proposed estimators are doubly robust, as they deliver valid inferences in the union model $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ where $\eta(x;\xi)$ is correctly specified, and either $p(z\mid u;\beta)$ or $p(y\mid u;\psi)$ is correctly specified. The proposed methodology requires correct specification of the model for $\eta(x)$, and therefore differs from the typical doubly robust inference when $\gamma=\gamma_{0}$ is known, where either the model for $\eta(x)$ or $p(y\mid R=1,x)$, but not necessarily both, is correctly specified. It also differs from the doubly robust inference with instrumental variables proposed by Sun et al., (2018) when $\gamma_{0}$ is unknown, where the model for $p(z\mid u)$ is correctly specified, and either the model for $\eta(x)$ or $p(y\mid R=1,x)$ is correctly specified. Specifying separate models for $\eta(x)$ and $p(y\mid R=1,x)$ in a way that respects exclusion restriction is difficult, as they are inextricably entwined in the conditional density $p(y\mid x)$. On the other hand, because $\psi$ and $\beta$ are variationally independent under exclusion restriction, the proposed methodology provides the analyst with two genuine independent opportunities to obtain valid inference.
@@ 2.4 Notation
Throughout, we use $A^{\mathrm{\scriptscriptstyle T}}$ to denote the transpose of a vector or matrix $A$, and $A_{1}\otimes A_{2}$ to denote the Kronecker product of two vectors or matrices $A_{1}$ and $A_{2}$. Estimation and inference are based on an independent and identically distributed observed data sample $(O_{1},...,O_{n})$, where $O=(R,RY,X)$. We denote the empirical measure as $\mathbb{P}_{n}$ so that empirical averages may be written as $n^{-1}\sum^{n}_{i=1}\{m(O_{i})\}=\mathbb{P}_{n}\{m(O)\}$. For an arbitrary function $f(y,x)$, denote $f^{{\dagger}}(y,x)=E\{f(Y,X)\mid X=x\}+E\{f(Y,X)\mid Y=y,U=u\}-E\{f(Y,X)\mid U=u\}$.
@@ 2.5 An example with binary outcome and instrumental variable
We illustrate the key ideas with binary outcome and instrumental variable, $\mathcal{Y}=\mathcal{Z}=\{0,1\}$. In this case, the exclusion restriction in assumption 2 is equivalent to $\textup{Cov}(Y,Z\mid U)=0$, which can be captured by the moment restriction $E[Y\{Z-p(Z=1\mid U)\}\mid U]=0$. The main idea is that we can replace the moment function with its inverse propensity weighted form, to create the following observed data conditional moment restriction for the tilting parameter $\gamma$,
$E\left[\frac{R}{\pi(Y,X;\eta,\gamma)}Y\{Z-p(Z=1\mid U)\}\biggr{\rvert}U\right]=0,$ (5)
@@
$E\left[\frac{R}{\pi(Y,X;\eta,\gamma)}{d}(U)Y\{Z-p(Z=1\mid U)\}\right]=0,$ (6)
where ${d}(u)\in\mathbb{R}^{p_{\gamma}}$ is a user-specified, vector-valued function with linearly independent elements. In principle, we can construct a semiparametric two-step estimator $\hat{\gamma}$ of $\gamma_{0}$ which solves the following empirical analogue of (6),
$\mathbb{P}_{n}\left[\frac{R}{\pi(Y,X;\hat{\eta},\gamma)}{d}(U)Y\{Z-\hat{p}(Z=1\mid U)\}\right]=0,$
where $\hat{\eta}(x)$ is a nonparametric estimator of $\eta(x)$ based on (2) for each fixed value of $\gamma$, and $\hat{p}(Z=1\mid u)$ is a nonparametric estimator of $p(Z=1\mid u)$. We can then construct the Hájek, (1971) estimator of the outcome mean,
$\hat{\mu}={\mathbb{P}_{n}\left\{\frac{RY}{\pi(Y,X;\hat{\eta},\hat{\gamma})}\right\}}\biggr{/}{\mathbb{P}_{n}\left\{\frac{R}{\pi(Y,X;\hat{\eta},\hat{\gamma})}\right\}}.$
On the other hand, the zero conditional covariance restriction under assumption 2 can also be captured by the alternative moment restriction $E[\{Y-p(Y=1\mid U)\}Z\mid U]=0$. This yields the observed data conditional moment restriction
$E\left[\frac{R}{\pi(Y,X;\eta,\gamma)}\{Y-p(Y=1\mid U)\}Z\biggr{\rvert}U\right]=0.$ (7)
@@
$E\left[\frac{R}{\pi(Y,X;\eta,\gamma)}\{Y-p(Y=1\mid U)\}\biggr{\rvert}U\right]=0.$ (8)
A key observation is that (7) and (8) are functionally different. We can then construct a semiparametric two-step estimator of $\gamma_{0}$ which solves the following empirical moment condition for some vector-valued function ${d}(u)\in\mathbb{R}^{p_{\gamma}}$,
$\mathbb{P}_{n}\left[\frac{R}{\pi(Y,X;\hat{\eta},\gamma)}{d}(U)\{Y-\hat{p}(Y=1\mid U)\}Z\right]=0,$
where $\hat{\eta}(x)$ is a nonparametric estimator of $\eta(x)$ based on (2), and $\hat{p}(Y=1\mid u)$ is a nonparametric estimator of $p(Y=1\mid u)$ based on (8), for each fixed value of $\gamma$. A Hájek, (1971) estimator of the outcome mean can be constructed similarly.
The preceding two approaches require estimation of $\eta(x)$, and either $p(z\mid u)$ or $p(y\mid u)$. In the presence of possibly high-dimensional $X$, we can perform inference under user-specified parametric specifications $\eta(x;\xi)$, $p(z\mid u;\beta)$ and $p(y\mid u;\psi)$. We obtain $\hat{\theta}=(\hat{\xi}^{\mathrm{\scriptscriptstyle T}},\hat{\beta}^{\mathrm{\scriptscriptstyle T}},\hat{\psi}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, for each fixed value of $\gamma$, as the solution to the following unconditional empirical moments
$\mathbb{P}_{n}\left[\left\{\frac{R}{\pi(Y,X;\xi,\gamma)}-1\right\}\frac{\partial\eta(X;\xi)}{\partial\xi}\right]=0,$
$\mathbb{P}_{n}\left\{{\frac{\partial\log p(Z\mid U;\beta)}{\partial\beta}}\right\}=0,$ (9)
$\mathbb{P}_{n}\left[\left\{\frac{R}{\pi(Y,X;\xi,\gamma)}\right\}{\frac{\partial\log p(Y\mid U;\psi)}{\partial\psi}}\right]=0.$
@@ 3.1 Semiparametric theory
The construction of the Hájek, (1971) estimators of $\mu_{0}$ in the previous section involves preliminary estimation of the tilting parameter $\gamma_{0}$. As $\gamma_{0}$ encodes the degree of departure from ignorability, it is also sometimes of interest in its own right in missing data problems. To simplify the presentation, in this section we consider joint estimation of the unknown $(p_{\gamma}+1)$-dimensional parameter ${{\phi}}_{0}=(\mu_{0},\gamma^{\mathrm{\scriptscriptstyle T}}_{0})^{\mathrm{\scriptscriptstyle T}}$. The construction of doubly robust estimators of ${{\phi}}_{0}$ is often motivated by the form of the influence function in $\mathcal{M}$ (Robins and Rotnitzky,, 2001; Chernozhukov et al.,, 2022). Specifically, any regular and asymptotically linear estimator $\hat{{\phi}}$ of ${{\phi}}_{0}$ in $\mathcal{M}$ satisfies
@@
where the $i$-th influence function $\Psi(O_{i})$ represents the the influence of the $i$-th observation on the estimator (Pfanzagl,, 1982; Bickel et al.,, 1993; Newey,, 1994; van der Laan and Robins,, 2003; Tsiatis,, 2007). The estimation theory for $\phi_{0}$ under general semiparametric models for the nonresponse mechanism has been previously developed (Robins et al.,, 2000; van der Laan and Robins,, 2003). The result below, proved in Sun et al., (2018), provides an application of the general theory to model $\mathcal{M}$.
@@
$\Psi(O;\phi_{0},c,d)=-\left[E\left\{\frac{\partial g(O;\phi,c,d)}{\partial\phi}\right\}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}g(O;\phi_{0},c,d),$
for some functions $c(y,x)\in\mathbb{R}$ and $d(y,x)\in\mathbb{R}^{p_{\gamma}}$, where
$g(O;\phi,c,d)=\frac{Rq(Y,X;\mu,c,d)}{\pi(Y,X;\gamma)}+{\left\{1-\frac{R}{\pi(Y,X;\gamma)}\right\}E\left\{q(Y,X;\mu,c,d)\rvert R=0,X;\gamma\right\}},$
@@
$q(Y,X;\mu,c,d)=\left\{\begin{matrix}Y-\mu+c(Y,X)-c^{{\dagger}}(Y,X)\\ d(Y,X)-d^{{\dagger}}(Y,X)\end{matrix}\right\}\in\mathbb{R}^{p_{\gamma}+1}.$
The function $g(O;\phi,c,d)$ has the familiar inverse probability weighted form in missing data literature, augmented by an additional term involving the nonrespondents (Robins et al.,, 1994, 2000; van der Laan and Robins,, 2003).
@@
In the special case with $\mathcal{Y}=\mathcal{Z}=\{0,1\}$, any function of the full data can be expressed as $f(Y,X)=f_{0}(U)+Yf_{y}(U)+Zf_{z}(U)+YZf_{yz}(U)$, where $\{f_{0}(\cdot),f_{y}(\cdot),f_{z}(\cdot),f_{yz}(\cdot)\}$ are arbitrary functions of the same dimension as $f(\cdot)$. It can be shown that
$f(Y,X)-f^{{\dagger}}(Y,X)$ $={f}_{yz}(U)\{Y-p(Y=1\mid U)\}\{Z-p(Z=1\mid U)\}.$
The example generalizes directly to discrete outcome and instrument taking values in $\mathcal{Y}=\{0,1,...,\ell_{y}\}$ and $\mathcal{Z}=\{0,1,...,\ell_{z}\}$, respectively. Let $v_{1}(Y)=\{I(Y=1),...,I(Y=\ell_{y})\}^{\mathrm{\scriptscriptstyle T}}$ and $v_{2}(Z)=\{I(Z=1),...,I(Z=\ell_{z})\}^{\mathrm{\scriptscriptstyle T}}$, where $I(\cdot)$ is the indicator function. Then for any function of the full data,
$f(Y,X)-f^{{\dagger}}(Y,X)={f}_{yz}(U)\{v_{1}(Y)-E(v_{1}(Y)\mid U)\}\otimes\{v_{2}(Z)-E(v_{2}(Z)\mid U)\},$

### v4

@@ 2.3 Alternative factorization
$p(r,w;\varphi)=\pi(w;\eta,\gamma)^{r}\{1-\pi(w;\eta,\gamma)\}^{1-r}\underbrace{p(y\mid u;\psi)p(z\mid u;\beta)p(u;\zeta)}_{p(w;\psi,\beta,\zeta)},$
where $\gamma$ is finite-dimensional and $(\eta,\psi,\beta,\zeta)$ is infinite-dimensional. The full data density is clearly separated from the nonresponse mechanism in the factorization above. This naturally encodes the exclusion restriction of assumption 2, which operates on the full data distribution, rather than on subpopulations defined by the nonresponse status. Furthermore, $\psi$ and $\beta$ are variationally independent, in the sense that any appropriate choice of $\psi$ and $\beta$ would result in a density in $\mathcal{M}$. Therefore, the models $p(y\mid u;\psi)$ and $p(z\mid u;\beta)$ can be specified separately without concerns about incompatibility. We show in the paper that estimation of $\mu_{0}$ requires consistent estimation of at least a subset of the nuisance parameters $(\eta,\psi,\beta)$. In modern studies, a broad collection of baseline covariates and operational features are usually recorded. When $X$ is high-dimensional or contains several continuous components, nonparametric estimation is typically infeasible in the moderate sample sizes that are found in practice, as the data are too sparse due to the curse of dimensionality (Robins and Ritov, 1997). Thus, we are often forced to specify more stringent dimension-reducing models $\eta(x;\xi)$, $p(y\mid u;\psi)$ and $p(z\mid u;\beta)$. Although in principle these models could be made as flexible as allowed by the sample size, we focus on parametric specifications in this paper to ease exposition.
@@
Accordingly, we define the submodels $\mathcal{M}_{1}$, $\mathcal{M}_{2}$ of $\mathcal{M}$ which correspond to models (1), (2) respectively. Thus, the proposed estimators are doubly robust, as they deliver valid inferences in the union model $\cup_{j=1,2}\mathcal{M}_{j}$ where $\eta(x;\xi)$ (and hence the parametric extended propensity score model) is correctly specified, and either $p(z\mid u;\beta)$ or $p(y\mid u;\psi)$ is correctly specified. The proposed methodology requires correct specification of the model for $\eta(x)$, and therefore differs from the typical doubly robust inference when $\gamma=\gamma_{0}$ is known, where either the model for $\eta(x)$ or $p(y\mid R=1,x)$, but not necessarily both, is correctly specified. It also differs from the doubly robust inference with instrumental variables proposed by Sun et al. (2018) when $\gamma_{0}$ is unknown, where the model for $p(z\mid u)$ is correctly specified, and either the model for $\eta(x)$ or $p(y\mid R=1,x)$ is correctly specified. Specifying separate models for $\eta(x)$ and $p(y\mid R=1,x)$ in a way that respects exclusion restriction is difficult, as they are inextricably entwined in the conditional density $p(y\mid x)$. On the other hand, because $\psi$ and $\beta$ are variationally independent under exclusion restriction, the proposed methodology provides the analyst with two genuine independent opportunities to obtain valid inference.
@@ 2.4 Notation
Throughout, we use $A^{\mathrm{\scriptscriptstyle T}}$ to denote the transpose of a vector or matrix $A$, and $A_{1}\otimes A_{2}$ to denote the Kronecker product of two vectors or matrices $A_{1}$ and $A_{2}$. Estimation and inference are based on an independent and identically distributed observed data sample $(O_{1},...,O_{n})$, where $O=(R,RY,X)$. We denote the empirical measure as $\mathbb{P}_{n}$ so that empirical averages may be written as $n^{-1}\sum^{n}_{i=1}\{m(O_{i})\}=\mathbb{P}_{n}\{m(O)\}$. For an arbitrary function $f(W)$ of the full data $W$, denote $f^{{\dagger}}(W)=\mathbb{E}\{f(W)\mid X\}+\mathbb{E}\{f(W)\mid Y,U\}-\mathbb{E}\{f(W)\mid U\}$.
@@ 2.5 An example with binary outcome and instrumental variable
We illustrate the key ideas with binary outcome and instrumental variable, $\mathcal{Y}=\mathcal{Z}=\{0,1\}$. In this case, the exclusion restriction in assumption 2 is equivalent to $\textup{Cov}(Y,Z\mid U)=0$, which can be captured by the moment restriction $\mathbb{E}[Y\{Z-p(Z=1\mid U)\}\mid U]=0$. The main idea is that we can replace the moment function with its inverse propensity weighted form, to create the following observed data conditional moment restriction for the tilting parameter $\gamma$,
$\mathbb{E}\left[\frac{R}{\pi(W;\eta,\gamma)}Y\{Z-p(Z=1\mid U)\}\biggr{\rvert}U\right]=0,$ (2.6)
@@
$\mathbb{E}\left[\frac{R}{\pi(W;\eta,\gamma)}{d}(U)Y\{Z-p(Z=1\mid U)\}\right]=0,$ (2.7)
where ${d}(u)\in\mathbb{R}^{p_{\gamma}}$ is a user-specified, vector-valued function with linearly independent elements. In principle, we can construct a semiparametric two-step estimator $\hat{\gamma}$ of $\gamma_{0}$ which solves the following empirical analogue of (2.7),
$\mathbb{P}_{n}\left[\frac{R}{\pi(W;\hat{\eta},\gamma)}{d}(U)Y\{Z-\hat{p}(Z=1\mid U)\}\right]=0,$
where $\hat{\eta}(x)$ is a nonparametric estimator of $\eta(x)$ based on (2.3) for each fixed value of $\gamma$, and $\hat{p}(Z=1\mid u)$ is a nonparametric estimator of $p(Z=1\mid u)$. We can then construct the following Hájek (1971) estimator of the outcome mean,
$\hat{\mu}={\mathbb{P}_{n}\left\{\frac{RY}{\pi(W;\hat{\eta},\hat{\gamma})}\right\}}\biggr{/}{\mathbb{P}_{n}\left\{\frac{R}{\pi(W;\hat{\eta},\hat{\gamma})}\right\}}.$
On the other hand, the zero conditional covariance restriction under assumption 2 can also be captured by the moment restriction $\mathbb{E}[\{Y-p(Y=1\mid U)\}Z\mid U]=0$. This yields the observed data conditional moment restriction
$\mathbb{E}\left[\frac{R}{\pi(W;\eta,\gamma)}\{Y-p(Y=1\mid U)\}Z\biggr{\rvert}U\right]=0.$ (2.8)
@@
$\mathbb{E}\left[\frac{R}{\pi(W;\eta,\gamma)}\{Y-p(Y=1\mid U)\}\biggr{\rvert}U\right]=0.$ (2.9)
A key observation is that (2.8) and (2.9) are functionally different. We can then construct a semiparametric two-step estimator of $\gamma_{0}$ which solves the following empirical moment condition for some vector-valued function ${d}(u)\in\mathbb{R}^{p_{\gamma}}$,
$\mathbb{P}_{n}\left[\frac{R}{\pi(W;\hat{\eta},\gamma)}{d}(U)\{Y-\hat{p}(Y=1\mid U)\}Z\right]=0,$
where $\hat{\eta}(x)$ is a nonparametric estimator of $\eta(x)$ based on (2.3), and $\hat{p}(Y=1\mid u)$ is a nonparametric estimator of $p(Y=1\mid u)$ based on (2.9), for each fixed value of $\gamma$. A Hájek (1971) estimator of the outcome mean can be constructed similarly.
The preceding two approaches require estimation of $\eta(x)$, and either $p(z\mid u)$ or $p(y\mid u)$. In the presence of possibly high-dimensional $X$, we can perform inference under working parametric specifications $\eta(x;\xi)$, $p(z\mid u;\beta)$ and $p(y\mid u;\psi)$. We obtain $\hat{\theta}=(\hat{\xi}^{\mathrm{\scriptscriptstyle T}},\hat{\beta}^{\mathrm{\scriptscriptstyle T}},\hat{\psi}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, for each fixed value of $\gamma$, as the solution to the following unconditional empirical moments,
$\mathbb{P}_{n}\left[\left\{\frac{R}{\pi(W;\xi,\gamma)}-1\right\}\frac{\partial\eta(X;\xi)}{\partial\xi}\right]=0;$
$\mathbb{P}_{n}\left\{{\frac{\partial\log p(Z\mid U;\beta)}{\partial\beta}}\right\}=0;$ (2.10)
$\mathbb{P}_{n}\left[\left\{\frac{R}{\pi(W;\xi,\gamma)}\right\}{\frac{\partial\log p(Y\mid U;\psi)}{\partial\psi}}\right]=0.$
@@ 3.1 Semiparametric theory
The construction of the Hájek (1971) estimators of $\mu_{0}$ in the previous section involves preliminary estimation of the tilting parameter $\gamma_{0}$. As $\gamma_{0}$ encodes the degree of departure from ignorability, it is also sometimes of interest in its own right in missing data problems. To simplify the presentation, in this section we consider joint estimation of the unknown $(p_{\gamma}+1)$-dimensional parameter ${{\phi}}_{0}=(\mu_{0},\gamma^{\mathrm{\scriptscriptstyle T}}_{0})^{\mathrm{\scriptscriptstyle T}}$. The construction of doubly robust estimators of ${{\phi}}_{0}$ is often motivated by the form of the influence function in $\mathcal{M}$ (Robins and Rotnitzky, 2001; Chernozhukov et al., 2022). Specifically, any regular and asymptotically linear estimator $\hat{{\phi}}$ of ${{\phi}}_{0}$ in $\mathcal{M}$ satisfies
@@
where the $i$-th influence function $\Psi(O_{i})$ represents the influence of the $i$-th observation on the estimator (Pfanzagl, 1982; Bickel et al., 1993; Newey, 1994; van der Laan and Robins, 2003; Tsiatis, 2007). The estimation theory for $\phi_{0}$ under general semiparametric models for the nonresponse mechanism has been previously developed (Robins et al., 2000; van der Laan and Robins, 2003). The result below, proved in Sun et al. (2018), provides an application of the general theory to model $\mathcal{M}$.
@@
$\Psi(O;\phi_{0},c,d)=-\left[\mathbb{E}\left\{\frac{\partial g(O;\phi,c,d)}{\partial\phi}\right\}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}g(O;\phi_{0},c,d),$
for some functions $c(W)\in\mathbb{R}$ and $d(W)\in\mathbb{R}^{p_{\gamma}}$ of the full data $W$, where
$g(O;\phi,c,d)=\frac{Rq(W;\mu,c,d)}{\pi(W;\gamma)}+{\left\{1-\frac{R}{\pi(W;\gamma)}\right\}\mathbb{E}\left\{q(W;\mu,c,d)\rvert R=0,X;\gamma\right\}},$
@@
$q(W;\mu,c,d)=\left\{\begin{matrix}Y-\mu+c(W)-c^{{\dagger}}(W)\\ d(W)-d^{{\dagger}}(W)\end{matrix}\right\}\in\mathbb{R}^{p_{\gamma}+1}.$
The function $g(O;\phi,c,d)$ has the familiar inverse probability weighted form in missing data literature, augmented by an additional term involving the nonrespondents (Robins et al., 1994, 2000; van der Laan and Robins, 2003).
@@
In the special case with $\mathcal{Y}=\mathcal{Z}=\{0,1\}$, any function of the full data $W$ can be expressed as $f(W)=f_{0}(U)+Yf_{y}(U)+Zf_{z}(U)+YZf_{yz}(U)$, where $\{f_{0}(U),f_{y}(U),f_{z}(U),f_{yz}(U)\}$ are functions of the same dimension as $f(W)$. It can be shown that
$f(W)-f^{{\dagger}}(W)$ $={f}_{yz}(U)\{Y-p(Y=1\mid U)\}\{Z-p(Z=1\mid U)\}.$
The example generalizes directly to discrete outcome and instrument taking values in $\mathcal{Y}=\{0,1,...,\ell_{y}\}$ and $\mathcal{Z}=\{0,1,...,\ell_{z}\}$, respectively. Let $v_{1}(Y)=\{I(Y=1),...,I(Y=\ell_{y})\}^{\mathrm{\scriptscriptstyle T}}$ and $v_{2}(Z)=\{I(Z=1),...,I(Z=\ell_{z})\}^{\mathrm{\scriptscriptstyle T}}$, where $I(\cdot)$ is the indicator function. Then for any function of the full data $W$,
$f(W)-f^{{\dagger}}(W)={f}_{yz}(U)\{v_{1}(Y)-\mathbb{E}(v_{1}(Y)\mid U)\}\otimes\{v_{2}(Z)-\mathbb{E}(v_{2}(Z)\mid U)\},$